A 3FT WIDE WALKWAY IS PLACED AROUND AN ANIMALS EXHIBIT AT THE ZOO. THE EXHIBIT IS RECTANGULAR IN SHAPE AND HAS LENGTH 15FT AND WIDTH 8 FT. WHAT IS THE AREA IN SQUARE FEET OF THE WALKWAY AROUND THE EXHIBIT.
Answer by Theo(13342)

the length of the exhibit is 15 feet.
the width of the exhibit is 8 feet.
those are also the dimensions of the inner edge of the walkway (the part that is right next to the exhibit).
the outer edge of the walkway has to be 3 feet more on each side.
this means the dimensions of the exhibit plus the walkway will be be 21 feet in length and 14 feet in width.
the area of the exhibit is equal to 15 * 8 = 120 square feet.
the area of the exhibit plus the walkway is equal to 21 * 14 = 294 square feet.
the difference between the area of the exhibit and the area of the exhibit plus the walkway is the area of the walkway.
that is 294 - 120 = 174 square feet.